OpenStudy (anonymous):

Help solve this question? 3/8x = x + 1/4

OpenStudy (anonymous):

deduct x from both sides: 3/8x-x =1/4 factor out x: x(3/8-1) = 1/4 solve.

OpenStudy (anonymous):

thanks for your help!